in reciprocating compressor with crank slider mechanism the obliquity ratio is 3.75 with crank radius of 0.2m the crank rotating at 20 rad/s what is the acceleration of the piston when the velocity is zero


1
Expert's answer
2021-11-01T08:51:38-0400

"A_p=w^2r[Cos\theta +\frac{Cos2\theta}{n}]"

"w=20rad\/s"

r=0.2m

"\theta =90 \because velocity=0"

n=3.75

"A_p=20^2*0.2[Cos90 +\frac{Cos2*90}{3.75}]"

=21.33m/s2